### v1

# EfficientQAT: Efficient Quantization-Aware Training for Large Language Models

Abstract
Large language models (LLMs) are integral to modern natural language processing and artificial intelligence. However, they face challenges in managing their significant memory requirements. Although quantization-aware training (QAT) offers a solution by reducing memory consumption through low-bit representations with minimal accuracy loss, it demands substantial training resources to optimize model weights and quantization parameters. To address this, we propose Efficient Quantization-Aware Training (EfficientQAT), a novel quantization technique for compressing LLMs. EfficientQAT involves two consecutive phases: Block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). Block-AP sequentially conducts quantization-aware training for all parameters in each transformer block with block-wise reconstruction, maintaining efficiency by avoiding training the entire LLM. Initialized with quantized model, E2E-QP then trains only quantization parameters (step sizes) end-to-end, enhancing efficiency with a fixed quantized backbone and reduced trainable parameter count. Extensive experiments demonstrate that EfficientQAT outperforms previous quantization methods across a range of models, including base LLMs, instruction-tuned LLMs, and multimodal LLMs, with scales from 7B to 70B parameters at various quantization bits. For instance, EfficientQAT obtains a 2-bit Llama-2-70B model on a single A100-80GB GPU in 41 hours, with less than 3% accuracy degradation compared to the full precision (69.48 vs. 72.41). Notably, this INT2 quantized 70B model obtains a 1.67 accuracy gain over the Llama-2-13B model (69.48 vs. 67.81) while requiring less memory (19.2GB vs. 24.2GB). Code is available at https://github.com/OpenGVLab/EfficientQAT.
1 Introduction
Recent advancements in large language models (LLMs) llama2 ; gpt4 ; vicuna ; lvlm-ehub ; mmtbench have demonstrated impressive capabilities in diverse language tasks such as reasoning arc ; boolq ; hellaswag , cognitive processing mme ; lvlm-ehub , and agent-based applications tool_learning ; toolllm . However, these models are characterized by their extensive parameters, which pose significant challenges for memory footprint and bandwidth squeezellm ; besa .
Quantization-aware training (QAT), one of the most effective quantization techniques, works by minimizing quantization errors through training with quantization constraints. Although QAT can compress LLMs effectively without significant performance loss, it requires training the whole LLM on a large corpus, resulting in enormous training costs. For instance, the QAT method BitNet b1.58 bitnet158 can achieve nearly lossless ternary quantization but requires retraining LLMs from scratch using the full pre-trained dataset, which is impractical for extremely large models***BitNet b1.58 only verifies the performance on 3B models with 100B training tokens..
In pursuit of efficient quantization for large language models (LLMs), techniques such as post-training quantization (PTQ) and quantized parameter-efficient fine-tuning (Q-PEFT) have been developed. PTQ awq ; gptq ; omniquant ; quip ; aqlm minimizes memory footprint during inference by converting pre-trained LLM weights from 16-bit to formats like 2-bit without retraining. While PTQ is suitable for quick deployment with minimal resources, it may compromise accuracy, especially in low-bit regimes, as shown in Figure 1(a). Q-PEFT qlora ; qa-lora ; loftq ; lq-lora ; peqa , on the other hand, uses parameter-efficient fine-tuning methods like LoRA lora to train quantized models, significantly reducing memory requirements during training. For instance, QLoRA qlora converts the model from FP16 to NF4 (a 4-bit float format) and then applies a LoRA module, allowing fine-tuning on consumer-grade GPUs without sacrificing performance. However, most Q-PEFT techniques qlora ; lq-lora ; loftq revert to FP16 for integrating additional LoRA modules post fine-tuning. This necessitates another round of PTQ for deployment on memory-limited platforms, which can degrade performance. Although solutions like PEQA peqa and QA-LoRA qa-lora attempt to address this issue, their performance remains suboptimal in low-bit scenarios (2-bit and 3-bit) due to poor quantized model initialization (Figure 1(b)).
(a) 2-bit quantization comparisons
(b) Q-PEFT comparisons
Figure 1: (a) EfficientQAT significantly surpasses existing uniform quantization methods, and is either superior to or comparable with vector quantization techniques. (b) EfficientQAT markedly outperforms existing Q-PEFT methods.
(a) inference speed comparisons
(b) training speed comparisons
Figure 2: (a): This panel shows a speed comparison of vector-matrix multiplications for the gate_proj component of Llama-2-70B. Uniform Quantization proves more hardware-efficient than vector quantization. (b): EfficientQAT can complet the QAT of 70B models on a single A100 GPU with less time.
To tackle the above challenges, we introduce a novel quantization technique named EfficientQAT, to compress LLMs effectively and efficiently. EfficientQAT surpasses existing PTQ omniquant ; aqlm , Q-PEFT qa-lora ; qlora and QAT llmqat methods in performance while maintaining memory efficiency during training and inference, as illustrated in Figure 2(a) & 2(b). To improve quantization efficiency and accuracy, EfficientQAT involves two consecutive phases: Block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). Block-AP sequentially conducts quantization-aware training for all parameters including the original full-precision weights and quantization parameters (step sizes and zero points) in each transformer block with block-wise reconstruction. While Block-AP trains all parameters, it preserves memory efficiency by avoiding training the entire LLM. Moreover, potential overfitting issues in Block-AP can be mitigated by simply increasing the training samples from 128 to 4096 (see Figure 5). Notably, our fully trainable paradigm of Block-AP outperforms other trainable variants (see Table 7), such as training rounding adaround ; weight_round , clipping thresholds omniquant , and step sizes lsq ; cbq .
Building on the robust initialization provided by Block-AP, E2E-QP fixes the quantized weights and trains only the quantization parameters (step sizes) end-to-end, further enhancing the quantization performance. Moreover, E2E-QP enables instruction tuning of quantized models for specific tasks in a quantization-aware and memory-efficient manner, as trainable quantization parameters take only a small fraction of the total network (e.g., 1.5% with a quantization group size of 64). It works like previous quantized parameter-efficient fine-tuning (Q-PEFT)techniques such as QA-LoRA qa-lora . Therefore, EfficientQAT is general enough to compress various models including base LLMs (Llama llama2 ; llama , instruction-tuned LLMs (Vicuna vicuna ), and multimodal LLMs (LLaVA llava1.5 ).
Thanks to the integration of the Block-AP and E2E-QP, EfficientQAT characterizes itself as a fast-converging, memory-efficient, and high-performing quantization technique. For instance, EfficientQAT can obtain a 2-bit Llama-2-70B model on a single A100-80GB GPU in just 41 hours, with less than 3% zero-shot accuracy degradation compared to its full-precision counterpart (69.48 vs. 72.41). We also evaluate EfficientQAT across scenarios involving model compression and instruction-tuning. In model compression, as illustrated in Figure 1(a), EfficientQAT significantly outperforms other uniform quantization methods by approximately 5 points on accuracy in the challenging 2-bit quantization setting. It also matches the performance of vector quantization methods aqlm ; quip-sharp in this scenario. Furthermore, Figure 2(a) demonstrates that the uniform quantization implemented by EfficientQAT is more hardware-efficient than vector quantization, which introduces a considerable overhead. In terms of instruction tuning, as shown in Figure 1(b), EfficientQAT consistently outperforms existing Q-PEFT methods, including QLoRA qlora , QA-LoRA qa-lora , and PEQA peqa .
2 Related Works
Post-Training Quantization of LLMs. PTQ is a pivotal technique for accelerating and deploying LLMs. Quantization approaches generally fall into two categories: weight-only quantization gptq ; spqr ; owq ; squeezellm ; norm_tweaking ; teq and weight-activation quantization smoothquant ; qllm ; os ; os+ ; rptq ; atom ; quik ; odyssey ; quarot . Weight-only quantization focuses on compressing weights into low-bit formats, reducing memory demands and enhancing the efficiency of memory-bounded computations in LLMs qserve ; llm-unveil . Conversely, weight-activation quantization compresses both weights and activations, thus further decreasing the overhead associated with matrix multiplications qserve . Recent advancements in weight-only quantization include the introduction of vector quantization methods by QUIP#quip-sharp and AQLMaqlm . These methods have shown promising performance but also introduce significant overhead llm-qbench . Our research continues to explore uniform quantization, which is preferred for its compatibility with hardware implementations.
Quantization-Aware Training of LLMs. QAT can enhance the performance of quantized models beyond what PTQ offers. However, QAT has been less explored in LLMs due to the significant training costs involved. Studies such as LLM-QAT llmqat and BitDistiller bitdistiller investigate the application of knowledge distillation within QAT contexts. Techniques like BitNet b1.58 bitnet158 and OneBit onebit employ QAT to achieve extreme binary or ternary quantization levels. Although BitNet b1.58 demonstrates near-lossless performance on models up to 3 billion parameters and 100 billion training tokens with ternary quantization, its applicability to larger models or datasets remains uncertain due to prohibitive training expenses.
Quantized Parameter-Efficient Fine-Tuning of LLMs. Techniques like QLoRA qlora , INT2.1 int2 , LQ-LoRA lq-lora , and LoftQ loftq quantize model parameters to low-bit representations followed by the addition of LoRA lora modules for fine-tuning. However, these methods require merging the LoRA modules into quantized weights, resulting in the model reverting to the FP16 format. Addressing this issue, QA-LoRA qa-lora redesign the LoRA module to merge seamlessly into the zero points. To our knowledge, the work most closely to our approach is PEQA peqa , which employs a simple round-to-nearest (RTN) method for low-bit quantization and fine-tunes the step sizes for downstream task adaptation. Despite this, the suboptimal PTQ initialization technique utilized by PEQA leads to performance that falls short when compared to QLoRA qlora and QA-LoRA qa-lora .
3 EfficientQAT
3.1 Method Overview
In this section, we introduce EfficientQAT, a novel quantization-aware training framework for LLMs that enhances memory efficiency. As illustrated in Figure 3, traditional QAT approaches train the full-precision weights $\mathbf{W}$ and quantization parameters $s$ (step sizes) and $z$ (zero points) simultaneously in an end-to-end manner, which significantly increases the memory requirements due to the large number of parameters involved. To address this issue, EfficientQAT adopts a two-stage strategy: block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). In the Block-AP phase, model parameters and quantization parameters are trained block-by-block using reconstruction loss, which not only allows for precise calibration with full training but also reduces memory consumption brecq ; omniquant by block-wise training. Following this, the E2E-QP phase fixes the quantized weights and trains the step sizes exclusively on target datasets, thus achieving quantization-aware training in a memory-efficient way. Details on Block-AP and E2E-QP are further described in Sections 3.2 and 3.3, respectively.
Figure 3: The overall pipeline of naive QAT and proposed EfficientQAT. EfficientQAT introduces two novel processes: Block-wise Training of All Parameters (Block-AP) and End-to-End Training of Quantization Parameters (E2E-QP).
3.2 Block-Wise Training of All Parameters
In this section, we introduce the Block-Wise Training of All Parameters (Block-AP) approach, designed to efficiently provide an effective initialization for following end-to-end training.
Quantization and Dequantization. Specifically, Block-AP begins with a standard uniform quantization method:
$\mathbf{W}_{int}=\mathrm{clamp}(\lfloor\frac{\mathbf{W}}{s}\rceil+z,0,2^{N}-1),$ (1)
[…]
$\widehat{\mathbf{W}}=(\mathbf{W_{int}}-z)\cdot s.$ (2)
Here, $\widehat{\mathbf{W}}$ refers to the dequantized weights used in the forward computation. The processes of quantization (Eq.(1)) and dequantization (Eq.(2)) are integrated within the computation graph and can be optimized through gradient descent in a quantization-aware manner.
Blcok-wise Quantization-aware Training. Traditional QAT methods bitnet158 ; lsq ; llmqat train the entire network using Eq.(1) and Eq.(2) in an end-to-end fashion, which typically requires substantial computational resources and extensive data to prevent overfitting. Here we aim to enhance the training efficiency of QAT. Previous studies, such as BRECQ brecq , have demonstrated that block-wise training achieves faster convergence and requires less training time, data, and memory than end-to-end training given a pre-trained model. Following the methodologies in BRECQ brecq and OmniQuant omniquant , Block-AP sequentially conducts quantization-aware training within one transformer block before moving on to the next under a block-wise reconstruction framework.
Full Training of Model Weights and Quantization Parameters. Unlike previous methods which optimize several quantization parameters such as rounding parameters adaround ; weight_round ; flexround , clipping parameters omniquant , and step sizes lsq ; cbq , Block-AP behaves like QAT, training all inherent parameters from Eq.(1) and Eq.(2), including scaling factor $s$, zero point $z$, and model weights $\mathbf{W}$.
In our Block-AP approach, a straightforward full-training regimen outperforms existing partial-training variants adaround ; brecq ; cbq with intricate designs. Traditional training methods involving rounding parameters adaround ; brecq ; cbq serve as regularization techniques, constraining the update range of integral weights to $(-1,+1)$ to mitigate overfitting. However, this approach limits the solution space, potentially hindering the final performance of quantized models. By simply scaling the training samples from $128$ to $4096$, Block-AP effectively addresses the overfitting issue (see Figure 5 for details) while improving performance compared to training with rounding parameters. Our empirical findings demonstrate the superiority of full training within our Block-AP over existing partial-training variants adaround ; brecq ; cbq , as shown in Table 7.
Following block-wise training, we obtain the quantized model which includes quantized weights $\mathbf{W}_{q}$, step sizes $s$, and zero points $z$ for each quantization group. The weights $\mathbf{W}_{q}$ and zero points $z$ are stored in a low-bit format, while step sizes $s$ are stored in FP16. Note that $s$ and $z$ are shared within their respective quantization groups and constitute only a small fraction of the model’s parameters, approximately 1.6% for a group size of 64. Moreover, the model’s memory footprint is substantially reduced by transitioning from full-precision 16-bit weights to 2/3/4-bit quantized weights.
3.3 End-to-End Training of Quantization Parameters
We further introduce the End-to-End Training of Quantization Parameters (E2E-QP), aimed at efficiently training the entire quantized model on target datasets.
End-to-End Training of step sizes. Unlike traditional Quantization-Aware Training (QAT) methods llmqat ; bitnet158 that train full-precision weights, E2E-QP begins with $\mathbf{W}_{q}$ initialized via Block-AP and focuses solely on the training of quantization parameters ($s$ and $z$). Our findings indicate that training $s$, $z$, or both yields similar performance (see Table 7 for details). However, since training $z$ involves converting it from a low-bits format to full-precision, we typically train only $s$ by default unless specified otherwise to avoid additional memory overhead.
Additionally, within E2E-QP, there is no quantization process as per Equation (1); only the dequantization process occurs as described in Equation (2). Thus, the gradient of the trainable parameter $s$ is computed as $\frac{\partial\widehat{w}}{\partial s}=w_{q}-z$.
Overall, the memory usage for training in E2E-QP is drastically reduced due to the reduced trainable parameter count. Detailed memory footprints for various model sizes and bits under E2E-QP are listed in Table 8. For instance, the Llama-2-70B model can complete 2-bit QAT through E2E-QP using only 34.2GB of memory. Equipped with E2E-QP, EfficientQAT is adaptable to different scenarios by simply changing the training datasets, which includes applications such as continual pre-training and instruction-tuning alpaca .
Table 1: Llama 2 & 3 average zero-shot accuracy on 5 common-sense reasoning tasks ($\uparrow$). Detailed number of each task can be found at Table 13 (4-bit), Table 14 (3-bit), and Table 15 (2-bit). “Group” indicates group size for uniform quantization and codebook scheme for vector quantization.
Method Bits Type Group (code) 2-7 2-13 2-70 3-8 3-70
FP16 16 - 16 64.86 67.81 72.41 68.58 75.33
RTN 4 uniform 128 64.52 67.50 72.26 67.79 73.98
GPTQ 4 uniform 128 64.24 67.27 72.39 67.80 74.74
AWQ 4 uniform 128 64.54 67.61 72.44 68.24 74.77
OmniQ 4 uniform 128 64.52 67.10 72.39 - -
AutoRound 4 uniform 128 64.39 67.36 72.47 - -
QuIP# 4 vector - 64.48 67.28 72.17 - -
EfficientQAT 4 uniform 128 64.27 67.52 72.62 68.43 74.57
RTN 3 uniform 128 62.06 65.77 70.83 58.72 65.29
GPTQ 3 uniform 128 62.48 66.18 71.47 60.58 71.28
AWQ 3 uniform 128 62.82 66.14 71.41 64.82 73.65
OmniQ 3 uniform 128 62.42 66.18 71.07 - -
AutoRound 3 uniform 128 63.72 66.68 71.24 - -
QuIP# 3 vector - 63.52 66.26 72.13 - -
EfficientQAT 3 uniform 128 64.02 67.28 71.76 67.35 72.42
OmniQ 2 uniform 128 46.98 53.56 54.87 - -
AutoRound 2 uniform 128 54.50 60.72 67.70 - -
EfficientQAT 2 uniform 128 59.50 63.88 68.93 59.37 67.57
AQLM 2 vector 2x8 57.61 62.22 69.85 - -
AQLM 2 vector 1x16 61.85 64.95 70.84 64.10 70.10
QuIP# 2 vector - 60.61 64.44 70.91 - -
EfficientQAT 2 uniform 64 60.14 64.48 69.48 60.76 67.89
4 Experiments
This section presents extensive experiments to verify our proposed EfficientQAT. Secition 4.1 and Sec 4.2 present the comparisons with quantization methods and Q-PEFT methods respectively. Section 4.4 details the training cost and inference speed-up of the proposed EfficientQAT. Section 4.3 presents the comprehensive ablation studies of the proposed EfficientQAT.
4.1 EfficientQAT for LLMs Quantization
[…]
Evaluation. We assess the zero-shot accuracy of five common-sense reasoning tasks using the v0.4.2 lm-evaluation-harnesshttps://github.com/EleutherAI/lm-evaluation-harness. The tasks include WinoGrande winogrande , PIQA piqa , HellaSwag hellaswag , Arc-Easy arc , and Arc-Challenge arc . We also measure the perplexity of Wikitext2 and C4 with a 2048 context length, as done in previous studies gptq ; omniquant .
PTQ Baseline. We compare our results with PTQ methods from uniform quantization such as GPTQ gptq , AWQ awq , OmniQ omniquant , and AutoRound weight_round , and vector quantization including QuIP# quip-sharp and AQLM aqlm .
Table 2: Comparison with QAT methods on Llama-2 & 3 models.
Model Method Bits Group PPL Average
Wiki C4 Accuracy
2-7B LLM-QAT 3 128 6.02 - -
2-7B BitDistiller 3 128 5.97 - -
2-7B EfficientQAT 3 128 5.81 7.34 64.02
2-7B LLM-QAT 2 128 9.30 - -
2-7B BitDistiller 2 128 8.08 - -
2-7B EfficientQAT 2 128 7.19 8.79 59.50
3-8B PB-LLM 2 128 24.70 79.20 38.80
3-8B DB-LLM 2 128 13.60 19.20 51.80
3-8B EfficientQAT 2 128 9.80 13.22 59.37
PTQ Results. The accuracy and perplexity results are presented in Table 1 and Table 3, respectively. It seems that minimal performance degradation with 4-bit group quantization, and even simple round-to-nearest (RTN) quantization achieves respectable results. Therefore, EfficientQAT offers a slight perplexity improvement ($\sim$0.02) in 4-bit uniform quantization and comparable zero-shot accuracy with existing methods. As the number of quantization bits decreases, EfficientQAT’s performance advantage increases. In 3-bit quantization, EfficientQAT surpasses the optimization-based method AutoRound by approximately 0.5% in zero-shot accuracy and outperforms OmniQuant by 0.14 to 0.43 in perplexity. Notably, in 2-bit quantization, EfficientQAT outperforms AutoRound weight_round by 5% in zero-shot accuracy for Llama-2-7b. EfficientQAT also exceeds previous vector quantization methods like AQLM aqlm using a 2x8 codebook. While AQLM aqlm with a 1x16 codebook and QuIP# quip-sharp perform slightly better in 2-bit scenarios, they introduce notable computational overhead llm-qbench and can even reduce inference speed, as shown in Figure 2(a). Thus, EfficientQAT successfully pushes the limits of uniform quantization in LLMs, making hardware-friendly uniform quantization competitive with more complex vector quantization methods. However, we can also find that Llama-3 models suffer more performance degeneration after quantization llama3_quant , which may caused by the full training with 15T tokens other than 2T tokens.
QAT Baseline. We also compare our results with existing QAT methods, including LLM-QAT llmqat , BitDistiller bitdistiller , PB-LLM pbllm and DB-LLM dbllm .
QAT Results. As demonstrated in Table 2, EfficientQAT markedly surpasses other QAT methods in performance. Specifically, EfficientQAT reduces perplexity by 0.89 compared to BitDistiller bitdistiller when applied to the Llama-2-7B model with w2g128 quantization. In the more challenging scenario of Llama-3 quantization llama3_quant , EfficientQAT exceeds the performance of DB-LLM dbllm by 5.98 in C4 perplexity and by 8.57 points in average accuracy with the w2g128 quantization.
Table 3: Llama 2 & 3 Wikitext2 and C4 perplexity ($\downarrow$), context length 2048.
Wikitext 2 C4
Method Bits Type Group 2-7 2-13 2-70 3-8 3-70 2-7 2-13 2-70 3-8 3-70
FP16 16 - 16 5.47 4.88 3.32 6.14 2.85 6.97 6.47 5.52 8.88 6.73
GPTQ 4 uniform 128 5.61 4.98 3.42 6.53 3.38 7.13 6.56 5.58 9.35 7.02
AWQ 4 uniform 128 5.62 4.97 3.41 6.55 3.26 7.13 6.56 5.58 9.41 6.98
OmniQ 4 uniform 128 5.58 4.95 3.40 - - 7.12 6.56 5.58 - -
QuIP# 4 vector - 5.56 4.95 3.38 - - 7.07 6.54 5.56 - -
EfficientQAT 4 uniform 128 5.53 4.93 3.39 6.47 3.17 7.07 6.54 5.58 9.26 6.94
GPTQ 3 uniform 128 6.29 5.42 3.85 9.58 5.25 7.89 7.00 5.85 11.66 8.64
AWQ 3 uniform 128 6.24 5.32 3.74 8.16 4.69 7.84 6.94 5.81 11.49 7.91
OmniQ 3 uniform 128 6.03 5.28 3.78 - - 7.75 6.98 5.85 - -
QuIP# 3 vector - 5.79 5.10 3.56 - - 7.32 6.72 5.67 - -
EfficientQAT 3 uniform 128 5.81 5.12 3.61 7.09 4.19 7.34 6.73 5.71 10.06 7.43
OmniQ 2 uniform 128 11.06 8.26 6.11 - - 15.02 11.05 8.52 - -
EfficientQAT 2 uniform 128 7.19 6.08 4.61 9.80 6.38 8.79 7.75 6.48 13.22 9.53
AQLM 2 vector 2x8 7.24 6.06 4.49 - - 8.96 7.80 6.36 - -
AQLM 2 vector 1x16 6.34 5.59 4.06 7.76 5.10 8.06 7.27 6.04 10.89 8.31
QuIP# 2 vector - 6.66 5.74 4.16 - - 8.35 7.45 6.12 - -
EfficientQAT 2 uniform 64 6.86 5.96 4.52 9.41 6.07 8.50 7.59 6.38 12.77 9.23
4.2 EfficientQAT for Instruction Tuning
Training and Evaluation. Following existing works qa-lora ; ir-qlora , we train Llama-1 models on the Alpaca dataset alpaca and assess their performance by measuring average 5-shot MMLU mmlu accuracy works qa-lora ; ir-qlora . The training hyperparameters are identical to those described in Section 4.1, except we replace the RedPajama dataset redpajama with Alpaca. In line with QLoRA’s methodology qlora , we adjust the source context length to 384 and the target context length to 128, training for 10,000 steps with a batch size of 16.
Baseline. We benchmark EfficientQAT against several leading methods, including QLoRA qlora , QA-LoRA qa-lora , PEQA peqa , and IR-QLoRA ir-qlora , across quantization setting of 2, 3, and 4 bits. Consistent with QA-LoRA qa-lora , we also employ GPTQ gptq to quantize the fine-tuned QLoRA models into a low-bit format without FP16 LoRA for equitable comparison.
Results. Both Table 7 and Figure 1(b) indicate that EfficientQAT significantly outperforms existing Q-PEFT methods. For instance, in channel-wise quantization (group size of -1), EfficientQAT achieves more than 3% higher accuracy than PEQA peqa . In the 2-bit quantization scenario, the superiority of EfficientQAT is even more pronounced, surpassing QA-LoRA qa-lora by 5.1% and 4.0% in 7B and 13B models, respectively, and outperforming PEQA by 4.5% and 8.7% in the same models. Moreover, Table 7 also demonstrates that EfficientQAT outperforms both QA-LoRA and QLoRA with GPTQ in smaller model memory footprint (larger group size).
Table 4: Llama-1 average MMLU accuracy (5-shot) about instruction-tuning on Alpaca dataset.
Method Bits Group 7B 13B
- 16 - 34.6 46.3
PEQA 4 -1 35.8 45.0
EfficientQAT 4 -1 38.8 48.2
QLoRA 4+16 - 38.4 48.4
QLoRA w/GPTQ 4 32 36.0 48.0
QA-LoRA 4 32 39.4 49.2
PEQA 4 64 39.4 47.4
IR-QLoRA 4 64 40.8 49.3
EfficientQAT 4 64 41.2 49.5
QLoRA w/ GPTQ 3 32 34.0 46.1
QA-LoRA 3 32 37.4 47.3
IR-QLoRA 3 64 38.4 -
PEQA 3 64 38.5 46.3
EfficientQAT 3 64 40.0 48.2
QLoRA w/ GPTQ 2 32 25.8 30.9
QA-LoRA 2 32 27.5 36.9
IR-QLoRA 2 64 27.8 -
PEQA 2 64 28.1 32.2
EfficientQAT 2 64 32.6 40.9
Table 5: Effectiveness of each component on Llama-2-7B w2g64 quantization.
Block-AP E2E-QP Avg. PPL Avg. Accuracy
453.49 40.69
8.53 58.99
9.33 55.71
7.68 60.14
Table 6: W2g64 Llama-2-7B performance with different trainable parameters in the block-wise training (w/o E2E-QP). “# Param.” indicates trainable parameters count in a block.
Param. # Param. Memory Avg. PPL Avg. Accuracy
clipping 6.3M 6.4GB 11.28 53.20
$s$,$z$ 6.3M 6.4GB 10.26 55.20
round 202.4M 8.6GB 15.50 45.32
$s$,$z$,round 208.7M 9.3GB 9.17 57.14
$s$,$z$,$\mathbf{W}$ 208.7M 8.5GB 8.65 58.94
Table 7: Llama-2-7B w2g64 quantization with different trainable parameters for E2E-QP (w/ Block-AP).
Param. Avg. Bits Avg. PPL Avg. Accuracy
$s$ 2.28 7.68 60.14
$z$ 2.50 7.69 60.08
$s,z$ 2.50 7.68 60.18
4.3 Ablation Analysis
The EfficientQAT algorithm is comprised of two main components: Block-AP and E2E-QP. This section evaluates the effectiveness, trainable parameters, and training sample requirements of each component. We present the average perplexity for WikiText2 and C4 datasets, and the average accuracy for five zero-shot reasoning tasks, similar to Table 1.
Effectiveness of each component. As indicated in Table 7, both the Block-AP and E2E-QP components significantly enhance performance, with their combination yielding the best results. Notably, Block-AP outperforms E2E-QP, aligning with findings from BRECQ brecq .
Trainable parameters of Block-AP. Block-AP trains all parameters, including original weights and quantization parameters. Previous methods have introduced various training strategies to mitigate overfitting, such as trained rounding adaround ; weight_round , clipping thresholds omniquant , and step sizes lsq ; cbq . We compare Block-AP with these methods by modifying only the trainable parameters of Block-AP. As shown in Table 7, Block-AP (training $s$, $z$, $\mathbf{W}$) performs best with an acceptable training cost. Additionally, the memory footprint of directly training $\mathbf{W}$ is even smaller than that of training the rounding operation, which requires an additional copy of rounding parameters.
Trainable parameters of E2E-QP. We further examine the trainable parameters within E2E-QP. Table 7 shows that training $s$, $z$, or both yields similar performance. However, given that converting $z$ from an original low-bit representation to a trainable FP16 format increases the average bit count, we opt to train only $s$ by default.
Samples number of Block-AP. We assess the number of training samples for Block-AP, noting that E2E-QP trains all parameters, which may lead to overfitting. To address this, we introduce an additional 64 unseen samples from ReadPajama to evaluate the overfitting issue. We adjust the training epochs to ensure a similar total training time, allowing for fair comparisons across different sample sizes. As illustrated in Figure 5, increasing the number of training samples significantly reduces the gap between training loss and validation loss from 1.07 to 0.06. This reduction corresponds to an increase in the average accuracy for zero-shot tasks from 57.14% to 58.99%. Consequently, we set the default number of training samples for E2E-QP at 4096, as this maintains a minimal gap between training and validation losses.
Samples number of E2E-QP. In the E2E-QP, we train the model for 1 epoch to avoid over-fitting. Our examination of the training sample sizes for E2E-QP, detailed in Table 5, reveals that average perplexity consistently improves as sample sizes increase from 128 to 32,674. However, there is no significant improvement in average accuracy beyond 4096 samples. Therefore, we set the training sample size for E2E-QP at 4096 by default to balance efficiency and performance. Nonetheless, it is possible to further enhance the performance of EfficientQAT by increasing the sample size.
Figure 4: Illustration of training loss, validation loss and average accuracy of w2g64 Llama-2-7b with different training samples size for Block-AP (w/o E2E-QP).
Figure 5: Llama-2-7B w2g64 quantization performance with different sample numbers for E2E-QP (w/ Block-AP).
# Samples Avg. PPL Avg. Accuracy
128 8.09 59.03
512 7.88 59.81
2048 7.75 60.13
4096 7.68 60.14
8192 7.63 60.19
32764 7.50 60.31
Table 8: The detailed training time and training memory of EfficientQAT across different model size and quantization bits on a single A100-80GB GPU.
Llama-2 Block-AP E2E-QP
Time Memory Time Memory (4-/3-/2-bits) Total Time
7B 3.3h 8.5GB $\sim$1.5h 7.0/6.4/5.6GB 4.8h
13B 5.6h 10.3GB $\sim$2.9h 11.7/10.6/9.1GB 8.5h
70B 26.6h 29.9GB $\sim$14.3h 48.4/42.0/34.2GB 40.9h
Table 9: Speed of the FP16 linear layer matrix-vector multiplication in PyTorch, and relative INT2 speedups in BitBLAS bitblas . Testing on A100-80GB GPU.
Llama-2 7B 13B 70B
size (out_c $\times$ in_c) 4096x4096 11008x4096 5120x5120 13824x5120 8192x8192 28672x8192
FP16 25 us 61 us 38 us 90 us 91 us 286 us
INT2 9 us 21 us 11 us 26 us 24 us 67 us
Speedup 3.1x 2.9x 3.6x 3.5x 3.9x 4.4x
4.4 Efficiency of EfficientQAT
Training Efficiency Table 8 illustrates the required memory and time for training Lllama-2 models using EfficientQAT. The results indicate that the model completes training rapidly, taking 4.8 hours for the 7B model and 40.9 hours for the 70B model. Furthermore, the adoption of a quantized backbone in the end-to-end quantization process (E2E-QP) significantly reduces memory usage due to the lower bit width. Consequently, EfficientQAT accomplishes 2-bit training for the 70B model using only 34.2GB of memory.
Inference Efficiency Weight-only quantization reduces the memory requirements of LLMs and accelerates inference, as LLMs are predominantly memory-bound llm-unveil . Quantization modifies the computation in the linear layers while other operations remain unchanged. To evaluate this, we calculate the relative speedup in the linear layers’ processing. According to Table 9, INT2 quantization enhances the forward-pass speed by approximately 2.9x to 4.4x. Furthermore, due to the leverage of standard uniform quantization, the quantized models of EfficientQAT can also achieve speedup through other toolbox, such as MLC-LLM mlc-llm , AWQ awq , and BitBLAS bitblas .
5 Conclusion
In this study, we introduce EfficientQAT, a novel method that completes QAT with improved efficiency in both memory usage and training time. Through comprehensive testing, EfficientQAT proves superior to existing PTQ, QAT, and Q-PEFT methods in terms of versatility and performance across various models and quantization levels. Additionally, EfficientQAT leverages a standard uniform quantization, which simplifies deployment using popular toolboxes. We anticipate that EfficientQAT will stimulate further research and improve the compression of Large Language Models (LLMs), making them more efficient and widely accessible.
Appendix A Limitations
In terms of limitations, EfficientQAT demonstrates marked superiority in low-bit scenarios, such as 3-bit and 2-bit quantization. However, in the 4-bit group-wise quantization, existing PTQ methods like GPTQ [22], AWQ [37], and even the simplest RTN achieve comparable performance more quickly. Additionally, the Llama-3 quantization experiences significant performance degradation [28], a phenomenon also observed in EfficientQAT. Moreover, the 2-bit performance of EfficientQAT still exhibits a small gap (3% to 4% accuracy) compared to full-precision models. Addressing these issues to achieve nearly lossless performance with INT2 quantization is a key focus of our future work.
Appendix B Broader Impact
This paper presents work whose goal is to advance the compression and acceleration of large language models. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix C Gradient of Trainable Parameters in Block-AP
Block-AP, aligned with LSQ+[4], uses a straight-through estimator (STE)[3] to facilitate gradient computation through the rounding operation. The gradients of scaling factor $s$ are computed as follows:
$\frac{\partial\widehat{w}}{\partial s}=\left\{\begin{aligned} &\lfloor\frac{w}{s}\rceil-\frac{w}{s},0\leq\lfloor\frac{w}{s}\rceil+z\leq 2^{N-1},\\ &-z,\lfloor\frac{w}{s}\rceil+z<0,\\ &2^{N-1}-z,\lfloor\frac{w}{s}\rceil+z>2^{N-1}.\end{aligned}\right.$ (3)
and the gradient with respect to zero point $z$ is:
$\frac{\partial\widehat{w}}{\partial z}=\left\{\begin{aligned} &0,0\leq\lfloor\frac{w}{s}\rceil+z\leq 2^{N-1},\\ &-1,otherwise,\\ \end{aligned}\right.$ (4)
[…]
$\frac{\partial\widehat{w}}{\partial w}=\left\{\begin{aligned} &1,0\leq\lfloor\frac{w}{s}\rceil+z\leq 2^{N-1},\\ &0,otherwise,\\ \end{aligned}\right.$ (5)
Appendix D Results Source of Other Method.
In this study, we present a thorough comparison of our method against existing PTQ techniques, including GPTQ [22], AWQ [37], OmniQ [52], AutoRound [11], QuIP# [58], and AQLM [20]. We also compare with existing QAT methods, including LLM-QAT [43], BitDistiller [19], PB-LLM [51] and DB-LLM [9]. Additionally, we also evaluate quantized parameter-efficient fine-tuning methods such as PEQA [29], QLoRA [16], QA-LoRA [64], and IR-QLoRA [47]. The results we discuss originate from their respective official publications, and other scholarly articles, or are derived from our reproduction. We meticulously document the source of the results for each method as follows:
  • GPTQ, AWQ, OmniQ, AutoRound: The zero-shot accuracy results for Llama-2 models using these methods are derived from the AutoRound GitHub repository§§§AutoRound: https://github.com/intel/auto-round/blob/main/docs/acc.md. The perplexity results for the Llama-2 models using GPTQ, AWQ, and OmniQ are taken from the OmniQ paper [52]. The results for Llama-3 models using AWQAWQ:https://github.com/mit-han-lab/llm-awq and GPTQGPTQ:https://github.com/qwopqwop200/GPTQ-for-LLaMa were obtained through their open-source implementations.
  • QuIP#, AQLM: We replicated the results using the official pre-trained models provided by QuIP#******https://github.com/Cornell-RelaxML/quip-sharp and AQLM††††††https://github.com/Vahe1994/AQLM.
  • LLM-QAT, BitDistiller: These results are cited from BitDistiller [19] paper.
  • PB-LLM, DB-LLM: These results are cited from recent Llama-3 quantization empirical study [28].
  • PEQA: The per-channel quantization results (g=-1) are cited from their publication [29], and the results for a group size of 64 were produced using our codebase.
  • QA-LoRA, QLoRA, QLoRA w/ GPTQ: These results are cited from QA-LoRA [64] paper.
  • IR-QLoRA: These results are cited from IR-QLoRA [47] paper.
Table 10: Model size of quantized models. Compression ratio indicates the compression ratio of quantized models compared with FP16 models.
Model # Bit Group size bits/param size (GiB) Compression ratio (%)
LLaMA-2-7B 16 - 16 12.55 -
[…]
4 64 4.31 3.74 70.20
4 128 4.16 3.62 71.14
[…]
3 64 3.30 3.13 75.08
3 128 3.15 3.01 75.98
[…]
2 64 2.28 2.21 82.40
2 128 2.14 2.10 83.25
LLaMA-2-13B 16 - 16 24.24 -
[…]
4 64 4.31 6.98 71.21
4 128 4.16 6.75 72.16
[…]
3 64 3.30 5.78 76.16
3 128 3.15 5.56 77.07
[…]
2 64 2.28 3.98 83.58
2 128 2.14 3.77 84.44
LLaMA-2-70B 16 - 16 128.48 -
[…]
4 64 4.31 35.34 72.49
4 128 4.16 34.10 73.46
[…]
3 64 3.30 28.87 77.53
3 128 3.15 27.67 78.46
[…]
2 64 2.28 19.16 85.09
2 128 2.14 18.04 85.96
Appendix E Size of Quantized Models
This section illustrates model size reduction achieved through quantization. Models quantized to low-bit representations are more compact.
We implement N-bit quantization with a grouping size of $g$, where each group of $g$ weights shares the same FP16 step size and an N-bit zero point. Consequently, the average number of bits per parameter is calculated as $N+\frac{N+16}{g}$. It is important to note that only the linear layers within the transformer blocks are quantized; other layers, such as normalization layers, embeddings, and the classification head, remain in FP16 format. Table 10 provides detailed comparisons of quantized model sizes and their compression ratios.
Table 11: Lllma-2-7B 2-bit quantization performance with different group sizes for proposed EfficientQAT.
Group Avg. Bits Avg. PPL Avg. Accuracy
32 2.56 7.59 60.28
64 2.28 7.68 60.14
128 2.10 7.99 59.50
256 2.07 8.18 58.67
Appendix F Additional Ablation Analysis
Quantization Group Size. The group size is a crucial hyperparameter in weight-only quantization. A smaller group size offers more granular compression and reduces quantization loss but increases the number of quantization parameters required. As indicated in Table 11, a group size of 64 strikes an optimal balance for 2-bit quantization using EfficientQAT. It outperforms a group size of 128 by achieving a 0.31 lower perplexity and a 0.64% higher accuracy, yet it slightly underperforms compared to a group size of 32, with a marginal difference of 0.09 in perplexity and 0.14% in accuracy.
Table 12: Results about instruction tuning of large vision-language models. We following the overall training pipeling of LLaVA-1.5 [39] and just change the fine-tuning methods. ‘QLoRA + Block-AP’ indicates that we leverage proposed Block-AP to quantized the QLoRA models into low-bits for fair comparisons. MME’s perception scores are normalized to 100 percent.
Model Method #Bit MMbench MME MM-Vet ScienceQA Avg.
Training Inference
LLaVA-1.5-7B LoRA 16 16 66.1 73.8 30.2 68.4 59.6
QLoRA 4+16 16 64.1 72.8 30.3 68.0 58.8
[…]
EfficientQAT 4 4 64.4 73.2 30.3 68.1 58.8(+0.5)
QLoRA + Block-AP 4+16 3 62.9 71.8 29.7 66.4 57.7
EfficientQAT 3 3 63.2 71.4 30.9 67.3 58.2(+0.5)
QLoRA + Block-AP 4+16 2 53.7 64.3 28.9 60.7 51.9
EfficientQAT 2 2 62.3 68.0 27.8 63.4 55.4(+3.5)
LLaVA-1.5-13B LoRA 16 16 68.5 77.1 38.3 71.2 63.8
QLoRA 4+16 16 67.6 76.9 36.0 69.9 62.7
[…]
EfficientQAT 4 4 67.5 74.8 35.6 70.2 62.0(-0.4)
QLoRA + Block-AP 4+16 3 66.8 75.5 34.5 68.4 61.3
EfficientQAT 3 3 67.4 74.8 35.3 69.3 61.7(+0.4)
QLoRA + Block-AP 4+16 2 62.5 72.1 32.5 65.0 58.0
EfficientQAT 2 2 63.9 73.1 33.9 68.6 59.9(+1.9)
Appendix G Instruction Tuning for LVLMs.
Traditional Q-PEFT methods only do experiments on the language models. In this section, we further extend proposed EfficientQAT into Large vision-Language models (LVLMs) such as LLaVA [40].
Training and Evaluation. For the fine-tuning of large vision-language models (LVLMs), we largely align with LLaVA1.5 [39], which encompass the training model, datasets, and hyperparameters‡‡‡‡‡‡For comprehensive details, please consult the official repository at https://github.com/haotian-liu/LLaVA.. Unlike LLaVA1.5, which begins fine-tuning with full-precision Vicuna models using either full fine-tuning or LoRA-based methods [27], EfficientQAT starts with Vicuna models already quantized using our Block-AP method and continues with our E2E-QP fine-tuning approach. The training process involves two steps: initially freezing the LLM and pre-training a projector to align features with a Vision Transformer (ViT), followed by end-to-end fine-tuning of both the LLM and the projector. For EfficientQAT, we modify the learning rates in the second step to $2\times 10^{-5}$ for 4-bit and $3\times 10^{-5}$ for 2-bit and 3-bit.
Evaluation. Evaluation of the fine-tuned LVLMs are conducted across four benchmarks: MME [23], MM-Vet [67], MMBench [42], and ScienceQA [44].
Baseline. We compare our results with those of QLoRA [16], applying our Block-AP method to quantize the QLoRA fine-tuned models to low bits for fair comparison.
Results. As shown in Table 12, EfficientQAT outperforms QLoRA [16] in low-bit settings for both LLaVA-1.5-7B and LLaVA-1.5-13B models, consistent with previous results in LMMs. Remarkably, the 2-bit LLaVA-1.5-13B model trained with EfficientQAT achieves an average score of 59.9, surpassing the 59.6 of the FP16 LLaVA-1.5-7B model trained with LoRA. However, there is a slight performance decrease observed in the 4-bit EfficientQAT and 16-bit QLoRA compared to the 16-bit LoRA, indicating that further research is needed to optimize Q-PEFT within LVLMs.
Appendix H Full Results
In Table 1, we present the average accuracy for five zero-shot tasks. This section offers a detailed breakdown of the task-specific accuracy numbers. Specifically, Tables 13, 14, and 15 detail the performance of 4-bit, 3-bit, and 2-bit quantization, respectively.
Table 13: 4-bit Llama 2 & 3 zero-shot accuracy by lm_eval v0.4.2 ( acc is reported, not acc_norm )
Model Method Bits Group WinoGrande HellaSwag ArcC ArcE PiQA Average $\uparrow$
2-7B - - 16 69.22 57.16 43.52 76.26 78.07 64.85
RTN 4 128 68.35 56.91 43.52 76.26 77.58 64.52
GPTQ 4 128 69.06 56.36 42.15 75.63 78.02 64.24
AWQ 4 128 68.98 56.40 43.86 76.14 77.31 64.54
OmniQ 4 128 68.98 56.59 43.34 75.76 77.91 64.52
AutoRound 4 128 68.67 56.79 42.58 75.76 78.13 64.39
QuIP# 4 - 69.22 56.65 43.00 75.51 78.02 64.48
EfficientQAT 4 128 69.22 57.00 42.32 75.13 77.69 64.27
2-13B - 16 - 72.22 60.07 48.29 79.42 79.05 67.81
RTN 4 128 72.14 59.77 47.95 79.00 78.62 67.50
GPTQ 4 128 72.14 59.76 47.53 78.58 78.35 67.27
AWQ 4 128 73.32 59.80 46.50 79.38 79.05 67.61
OmniQ 4 128 72.06 59.53 47.18 78.37 78.35 67.10
AutoRound 4 128 71.67 59.87 47.01 79.25 79.00 67.36
QuIP# 4 - 72.69 59.49 46.59 78.62 79.00 67.28
EfficientQAT 4 128 71.98 59.87 47.53 79.34 78.89 67.52
2-70B - 16 - 77.98 64.77 54.44 82.70 82.15 72.41
RTN 4 128 78.14 63.93 54.78 82.79 81.66 72.26
GPTQ 4 128 78.22 64.45 54.61 82.79 81.88 72.39
AWQ 4 128 77.58 64.48 55.12 82.70 82.32 72.44
OmniQ 4 128 77.51 64.52 55.12 82.91 81.88 72.39
AutoRound 4 128 78.30 64.60 54.52 82.87 82.05 72.47
QuIP# 4 - 77.82 64.51 54.44 82.37 81.72 72.17
EfficientQAT 4 128 78.45 64.57 55.12 83.00 81.94 72.62
3-8B - - 16 72.61 60.17 50.43 80.09 79.60 68.58
RTN 4 128 73.16 59.04 48.38 79.25 79.11 67.79
GPTQ 4 128 73.72 59.17 47.78 79.38 78.94 67.80
AWQ 4 128 73.01 59.43 50.00 79.55 79.22 68.24
EfficientQAT 4 128 72.53 59.43 50.94 79.84 79.43 68.43
3-70B - 16 80.51 66.36 60.41 86.99 82.37 75.33
RTN 4 128 78.77 65.83 57.34 85.69 82.26 73.98
GPTQ 4 128 80.51 66.12 59.04 85.77 82.26 74.74
AWQ 4 128 80.35 65.82 59.13 86.41 82.15 74.77
EfficientQAT 4 128 79.24 66.27 59.13 85.86 82.37 74.57
Table 14: 3-bit Llama 2 & 3 zero-shot accuracy by lm_eval v0.4.2 ( acc is reported, not acc_norm )
Model Method Bits Group WinoGrande HellaSwag ArcC ArcE PiQA Average $\uparrow$
2-7B - - 16 69.22 57.16 43.52 76.26 78.07 64.85
RTN 3 128 67.56 54.90 38.57 72.98 76.28 62.06
GPTQ 3 128 68.59 53.66 40.19 73.74 76.01 62.44
AWQ 3 128 67.40 54.98 41.64 74.07 76.01 62.82
OmniQ 3 128 66.69 54.42 39.85 74.37 76.77 62.42
AutoRound 3 128 68.27 55.33 42.92 75.25 76.82 63.72
QuIP# 3 - 68.19 55.85 41.89 74.62 77.04 63.52
EfficientQAT 3 128 69.14 55.90 42.83 74.66 77.58 64.02
2-13B - 16 - 72.22 60.07 48.29 79.42 79.05 67.81
RTN 3 128 70.72 57.74 44.62 77.69 78.07 65.77
GPTQ 3 128 70.88 57.83 45.65 77.99 78.56 66.18
AWQ 3 128 71.82 58.58 44.62 77.95 77.75 66.14
OmniQ 3 128 70.01 58.46 46.16 77.86 78.40 66.18
AutoRound 3 128 71.59 59.11 45.82 78.58 78.29 66.68
QuIP# - 3 72.45 58.26 44.62 77.90 78.07 66.26
EfficientQAT 3 128 72.06 59.01 47.95 79.00 78.40 67.28
2-70B - 16 - 77.98 64.77 54.44 82.70 82.15 72.41
RTN 3 128 77.90 61.98 52.39 81.10 80.79 70.83
GPTQ 3 128 77.66 62.94 53.67 81.65 81.45 71.47
AWQ 3 128 76.48 63.75 53.67 81.40 81.77 71.41
OmniQ 3 128 76.48 63.54 52.82 81.02 81.50 71.07
AutoRound 3 128 76.56 63.83 52.56 81.73 81.50 71.24
QuIP# 3 - 76.24 64.22 55.89 82.11 82.21 72.13
EfficientQAT 3 128 77.27 64.20 53.75 81.73 81.83 71.76
3-8B - - 16 72.61 60.17 50.43 80.09 79.60 68.58
RTN 3 128 66.54 50.87 36.69 65.36 74.16 58.72
GPTQ 3 128 70.88 55.13 37.80 65.24 73.83 60.58
AWQ 3 128 70.96 55.43 44.20 75.84 77.69 64.82
EfficientQAT 3 128 71.51 57.81 48.81 80.01 78.63 67.35
3-70B - 16 80.51 66.36 60.41 86.99 82.37 75.33
RTN 3 128 65.90 54.22 48.46 78.83 79.05 65.29
GPTQ 3 128 78.14 62.58 52.99 82.07 80.63 71.28
AWQ 3 128 78.85 64.26 58.36 84.51 82.26 73.65
EfficientQAT 3 128 77.82 65.53 55.12 83.12 80.52 72.42
Table 15: 2-bit Llama 2 & 3 zero-shot accuracy by lm_eval v0.4.2 ( acc is reported, not acc_norm )
Model Method Bits Group WinoGrande HellaSwag ArcC ArcE PiQA Average $\uparrow$
2-7B - - 16 69.22 57.16 43.52 76.26 78.07 64.85
GPTQ 2 128 55.17 32.59 21.25 40.45 58.32 41.56
OmniQ 2 128 55.88 40.28 23.46 50.13 65.13 46.98
AutoRound 2 128 61.01 40.28 32.25 65.99 72.96 54.50
AQLM 2 2x8 65.27 49.96 32.85 66.92 73.07 57.61
AQLM 2 1x16 65.19 53.42 39.68 74.07 76.88 61.85
QuIP# 2 - 65.67 52.19 37.88 71.84 75.46 60.61
EfficientQAT 2 128 66.22 50.84 36.52 69.78 74.16 59.50
EfficientQAT 2 64 65.98 51.58 36.86 70.96 75.30 60.14
2-13B - 16 - 72.22 60.07 48.29 79.42 79.05 67.81
GPTQ 2 128 55.80 41.06 21.93 55.60 67.08 48.29
OmniQ 2 128 57.93 46.23 30.29 63.22 70.13 53.56
AutoRound 2 128 64.33 53.35 38.57 71.17 76.17 60.72
AQLM 2 2x8 66.22 54.62 40.10 73.06 77.09 62.22
AQLM 2 1x16 70.09 57.62 43.52 75.25 78.29 64.95
QuIP# 2 - 69.06 56.53 42.92 75.72 77.97 64.44
EfficientQAT 2 128 68.90 55.66 42.83 75.04 76.99 63.88
EfficientQAT 2 64 68.36 55.27 41.89 74.83 77.04 63.48
2-70B - 16 - 77.98 64.77 54.44 82.70 82.15 72.41
GPTQ 2 128 49.57 25.04 22.70 25.08 49.51 34.38
OmniQ 2 128 64.33 35.45 33.28 67.21 74.10 54.87
AutoRound 2 128 74.90 59.65 46.59 78.37 79.00 67.70
AQLM 2 2x8 75.61 61.94 51.45 79.76 80.47 69.85
AQLM 2 1x16 76.01 62.78 52.99 81.36 81.07 70.84
QuIP# 2 - 75.77 62.86 52.65 81.90 81.39 70.91
EfficientQAT 2 128 73.64 61.58 49.23 80.01 80.20 68.93
EfficientQAT 2 64 74.59 61.78 50.77 80.13 80.14 69.48
3-8B - - 16 72.61 60.17 50.43 80.09 79.60 68.58
AQLM 2 1x16 71.82 55.44 41.21 74.24 77.80 64.10
EfficientQAT 2 128 65.67 50.74 36.01 69.15 75.30 59.37
EfficientQAT 2 64 67.72 51.86 37.03 71.17 76.03 60.76
3-70B - 16 80.51 66.36 60.41 86.99 82.37 75.33
AQLM 2 1x16 78.22 63.47 50.34 78.83 79.65 70.10
EfficientQAT 2 128 69.46 60.75 48.81 79.25 79.60 67.57
EfficientQAT 2 64 74.03 61.60 49.06 77.40 77.37 67.89